# MANet: Fine-Tuning Segment Anything Model for Multimodal Remote Sensing Semantic Segmentation

### v1

@@ Abstract
Multimodal remote sensing data, collected from a variety of sensors, provide a comprehensive and integrated perspective of the Earth’s surface. By employing multimodal fusion techniques, semantic segmentation offers more detailed insights into geographic scenes compared to single-modality approaches. Building upon recent advancements in vision foundation models, particularly the Segment Anything Model (SAM), this study introduces a novel Multimodal Adapter-based Network (MANet) for multimodal remote sensing semantic segmentation. At the core of this approach is the development of a Multimodal Adapter (MMAdapter), which fine-tunes SAM’s image encoder to effectively leverage the model’s general knowledge for multimodal data. In addition, a pyramid-based Deep Fusion Module (DFM) is incorporated to further integrate high-level geographic features across multiple scales before decoding. This work not only introduces a novel network for multimodal fusion, but also demonstrates, for the first time, SAM’s powerful generalization capabilities with Digital Surface Model (DSM) data. Experimental results on two well-established fine-resolution multimodal remote sensing datasets, ISPRS Vaihingen and ISPRS Potsdam, confirm that the proposed MANet significantly surpasses current models in the task of multimodal semantic segmentation. The source code for this work will be accessible at https://github.com/sstary/SSRS.
@@
Multimodal fusion, Remote Sensing, Semantic Segmentation, Segment Anything Model, Adapter
@@ I Introduction
Multimodal remote sensing semantic segmentation involves the process of classifying each pixel in remote sensing images using data from multiple sources or modalities, such as optical images, multispectral images, hyperspectral images, and LiDAR. By integrating diverse types of information, multimodal approaches enhance the accuracy and robustness of segmentation, particularly in complex environments [1, 2]. This technique leverages the complementary strengths of different data types to improve the identification of land objects, making it crucial for applications like land use and land cover [3, 4], environmental monitoring [5], and disaster management [6, 7]. In recent years, deep learning technologies have greatly prompted the development of multimodal fusion methods in remote sensing.
@@
Initially, convolution neural networks (CNNs) were the dominant architecture, known for their ability to extract local spatial features from different modalities with the encoder-decoder framework [8]. These early CNN-based methods to stack multimodal data and fuse features at various stages for improved segmentation performance [9, 10]. With the introduction of the self-attention-based Transformer [11], which excel at modeling global context and long-range dependencies, many hybrid methods integrating Transformers into CNN-based methods have emerged. In particular, Vision Transformer (ViT) [12] and Swin Transformer [13] further introduced Transformer into the field of computer vision, greatly improving the ability of image feature extraction. Combining CNNs for detailed feature extraction with Transformers for capturing global relationships between different data sources marked a new stage in segmentation models [14, 15, 16]. This hybrid approach enhances the ability to fuse information across modalities and scales, leading to more accurate, robust segmentation results in complex scenes. Despite their good performance, the aforementioned models were trained solely on task-specific data, meaning they lack the broad, general knowledge that models like foundation models possess, which were trained on vast amounts of diverse data beyond the scope of a single task.
Figure 1: Schematic of the SAM framework [17]. SAM is made up of three components, with its core image encoder built on ViT blocks.
@@
Interestingly, the recent emergence of large foundation models offers a solution to this challenge. In particular, the Segment Anything Model (SAM) [17] is a cutting-edge image segmentation model designed to tackle a wide range of segmentation tasks across diverse datasets. As presented in Fig. 1, SAM is comprised of three components, including a ViT-based image encoder, a prompt encoder and a mask decoder. Developed by Meta AI, SAM benefits from large-scale training on a vast visual corpus of natural images, enabling it to generalize effectively to unseen objects. Its versatility and robust performance position it as a valuable tool for various applications [18, 19]. However, compared to natural images, remote sensing images exhibit significant differences due to variations in sensors, resolution, spectral ranges, etc [20, 21, 22]. For multimodal tasks, non-optical data such as DSM further increases the discrepancy. This raises a crucial research problem: How can the capabilities of SAM, acquired from massive natural visual corpora, be leveraged to enhance multimodal remote sensing tasks?
@@
Existing fine-tuning techniques, such as Adapter [23, 24, 25] and LoRA [26], address part of this challenge. Compared to training and full fine-tuning, these methods keep most of the parameters frozen and learns task-specific information with very fewer parameters. This enables parameter-efficient learning and the successful migration of large foundation models to a wider range of specific downstream tasks, even in more constrained hardware environments. The concept of Adapter was first introduced in the natural language processing community [23] as a method to fine-tune large pre-trained models for specific downstream tasks. The core idea behind Adapter is to introduce a parallel, compact, and scalable adaptation module that learns task-specific knowledge during training while the original model branch remains fixed, retaining task-agnostic knowledge. Then the synergy of task-specific and task-agnostic knowledge facilitates efficient fine-tuning of the large model.
@@
In the field of remote sensing, fine-tuning techniques have focused on adapting SAM for single-modality tasks [27, 28]. CWSAM [27] and MeSAM [28] adapted SAM’s image encoder and introduced custom mask decoders for remote sensing data. However, as shown in in Table I, an analysis of the SAM parameter distribution across three scales reveals that the majority of SAM’s parameters are concentrated in its image encoder, suggesting that most of its general knowledge is encapsulated within this component. Therefore, we consider it unnecessary to modify SAM’s prompt encoder and mask decoder for remote sensing-specific tasks, as such modifications increase model adaptation complexity while hindering seamless integration with existing segmentation models.
TABLE I: The number of parameters in each part of the SAM under different backbones.
Backbone Image Encoder Prompt Encoder Mask Decoder
ViT-B 89.7M 6.2K 4.0M
ViT-L 308.3M 6.2K 4.0M
ViT-H 637.0M 6.2K 4.0M
@@
To address the aforementioned challenges, we introduce a modified multimodal adapter, namely MMAdapter for multimodal fusion remote sensing tasks. Specifically, it consists of two fine-tuned branches that learn modality-specific information and achieve adaptive multimodal fusion with two weighting factors. Leveraging the MMAdapter, we propose a multimodal fusion framework, namely multimodal adapter-based network (MANet). The MANet utilizes a Deep Fusion Module (DFM) to perform multiscale processing and fusion of the deep high-level features from the SAM’s image encoder. This is an effective solution for the complex characteristics of remote sensing images. Additionally, we employ a universal semantic segmentation decoder [29] that does not require additional task-specific design efforts, allowing for easy integration with decoders from other remote sensing tasks. The contributions of this work are threefold, as summarized in the following:
  • A novel multimodal adapter, namely MMAdapter, is proposed to extract and fuse multimodal remote sensing features by fine-tuning SAM with few parameters, which reveals the powerful generalization capability of SAM on DSM for the first time.
@@
  • A novel multimodal fusion network, namely MANet, is proposed by capitalizing on SAM’s image encoder and MMAdapter to perform remote sensing semantic segmentation. This is a streamlined and flexible adaptation framework that eliminates most of the redundant modules in SAM.
@@
  • Extensive experiments on two well-known fine-resolution multimodal remote sensing datasets, ISPRS Vaihingen and ISPRS Potsdam, confirm that the proposed MANet substantially outperforms existing methods.
@@
The remainder of this paper is structured as follows. Sec. II provides a review of related works on multimodal fusion and SAM. Sec. III presents the structure of the proposed MANet, followed by a detailed description of the extensive experiments in Sec. IV. Finally, the conclusion is given in Sec. V.
@@ II-A Multimodal Remote Sensing Semantic Segmentation
Semantic segmentation is a critical preprocessing step in remote sensing image processing, and leveraging multimodal information often yields better results than relying on a single modality. In recent years, the advent of deep learning has revolutionized the entire field of remote sensing, including semantic segmentation. Based on the classical encoder-decoder framework [8], numerous multimodal fusion approaches based on CNNs and Transformers have driven significant advancements in the field [30, 14, 15, 16]. ResUNet-a [31], an early architecture that based on CNNs, simply stacked multimodal data into four channels. vFuseNet [30] introduced a two-branch encoder to separately extract multimodal features, enabling deeper multimodal fusion through element-wise operations at the feature level. The introduction of Transformers [11, 12] has further enriched multimodal networks. For instance, CMFNet [14] employed CNNs for feature extraction and uses a Transformer structure to connect multimodal features across scales, emphasizing the importance of scale in multimodal fusion. Similarly, MFTransNet [16] used CNNs for feature extraction, while enhanced the self-attention module with spatial and channel attention for better feature fusion. FTransUNet [16] presented a multilevel fusion approach to refine the fusion of shallow- and deep-level remote sensing semantic features. Despite the great performance they achieved, we argue that existing models lack sufficient general knowledge, which poses a fundamental limitation to the advancement of multimodal fusion methods.
@@ II-B SAM in Remote Sensing
The Segment Anything Model (SAM) [17] holds a unique position as a general-purpose image segmentation model. This vision foundation model was trained on a very large visual corpus. It endows SAM with a remarkable ability to generalize to unseen objects, making it well-suited for applications in diverse scenarios. Nowadays, SAM has already been applied across various fields, such as autonomous driving [32], medical image processing [33] and remote sensing [34, 35, 36, 37]. In remote sensing, SAMRS [34] leveraged SAM to integrate numerous existing remote sensing datasets, using a new prompt named rotated bounding box. Furthermore, some recent works have considered fine-tuning SAM for remote sensing tasks such as semantic segmentation [27, 28, 38] and change detection [39, 40].
For single modality tasks, CWSAM [27] adapted SAM to synthetic aperture radar (SAR) images by introducing a task-specific input module and a class-wise mask decoder. MeSAM [28] incorporated an inception mixer into the SAM’s image encoder to retain high-frequency features and introduced a multiscale-connected mask decoder for optical images. For multimodal tasks, RingMo-SAM [38] introduced a prompt encoder tailored to the characteristics of multimodal remote sensing data, along with a category-decoupling mask decoder. It is observed that these methods focus on designing task-specific prompts and mask decoders. However, as discussed in Sec. I, the general knowledge in SAM is primarily centered in the image encoder. While these approaches successfully utilize SAM’s general knowledge to remote sensing, their complex architectures greatly hinder their adaptation to existing general semantic segmentation networks. Furthermore, there is currently no SAM-based multimodal approach designed for DSM data.
@@ III Proposed Method
To begin with, we first present the conventional single-modal Adapter in Sec. III-A before introducing the proposed MMAdapter in Sec. III-B. Building upon the proposed MMAdapter, the proposed MANet is elaborated in Sec. III-B.
III-A The Standard Adapter
Figure 2: (a) The ViT block without Adapter in SAM’s image encoder and (b) the ViT block equipped with standard Adapter [41]. The Adapter facilitates the efficient utilization of general knowledge in specific tasks.
Figure 3: (a) The ViT block endowed with the proposed MMAdapter and (b) the detailed structure of the MMAdapter. Compared to the standard Adapter, MMAadapter is characterized by two branches of shared weights for multimodal feature extraction. The standard Adapter and the proposed MMAdapter are used to fine-tune and fuse features, respectively.
Figure 4: Overview of the proposed MANet consisting of a SAM’s image encoder with MMAdapter, a DFM and a general Decoder.
As discussed in Sec.I, SAM general knowledge is confined to its image encoder, specifically the ViT block, whose structure is shown in Fig.2(a). In [41], the Adapter was proposed to enhance the capabilities of ViT blocks for medical tasks through fine-tuning, as illustrated in Fig. 2(b). Instead of adjusting all parameters, the pre-trained SAM parameters remain frozen, while two Adapter modules are introduced to learn task-specific knowledge. Each Adapter consists of a down projection, a ReLU activation, and an up projection. The down projection compresses the input embedding to a lower dimension using a simple MLP layer, and the up projection restores the compressed embedding to its original dimension using another MLP layer. For a specific input feature $\bm{x}_{i}\in\mathbb{R}^{h\times w\times c}$, where $h$, $w$, and $c$ represent the height, width, and channels of the input feature, the Adapter’s process for generating the adapted feature can be expressed as:
$\bm{x}_{a}$ $=$ $\mathrm{ReLU}\left(\mathrm{LN}(\bm{x}_{i})\cdot\bm{W}_{d}\right)\cdot\bm{W}_{u},$ (1)

### v2

# A Unified Framework with Multimodal Fine-tuning for Remote Sensing Semantic Segmentation

@@ Abstract
Multimodal remote sensing data, acquired from diverse sensors, offer a comprehensive and integrated perspective of the Earth’s surface. Leveraging multimodal fusion techniques, semantic segmentation enables detailed and accurate analysis of geographic scenes, surpassing single-modality approaches. Building on advancements in vision foundation models, particularly the Segment Anything Model (SAM), this study proposes a unified framework incorporating a novel Multimodal Fine-tuning Network (MFNet) for remote sensing semantic segmentation. The proposed framework is designed to seamlessly integrate with various fine-tuning mechanisms, demonstrated through the inclusion of Adapter and Low-Rank Adaptation (LoRA) as representative examples. This extensibility ensures the framework’s adaptability to other emerging fine-tuning strategies, allowing models to retain SAM’s general knowledge while effectively leveraging multimodal data. Additionally, a pyramid-based Deep Fusion Module (DFM) is introduced to integrate high-level geographic features across multiple scales, enhancing feature representation prior to decoding. This work also highlights SAM’s robust generalization capabilities with Digital Surface Model (DSM) data, a novel application. Extensive experiments on three benchmark multimodal remote sensing datasets, ISPRS Vaihingen, ISPRS Potsdam and MMHunan, demonstrate that the proposed MFNet significantly outperforms existing methods in multimodal semantic segmentation, setting a new standard in the field while offering a versatile foundation for future research and applications. The source code for this work is accessible at https://github.com/sstary/SSRS.
@@
Multimodal Fine-tuning, Remote Sensing, Semantic Segmentation, SAM, Adapter, LoRA
@@ I Introduction
Multimodal remote sensing semantic segmentation involves the process of classifying each pixel in ground images using data from multiple sources or modalities, such as optical images, multispectral images, hyperspectral images, and LiDAR. By integrating diverse types of information, multimodal approaches enhance the accuracy and robustness of segmentation, particularly in complex environments [gomez2015multimodal, li2022deep]. This technique leverages the complementary strengths of different data types to improve the identification of land objects, making it crucial for applications such as land use and land cover [yao2023extended, hong2021multimodal], environmental monitoring [karmakar2023crop], and disaster management [algiriyage2022multi, zhang2023cross]. In recent years, deep learning technologies have greatly prompted the development of multimodal fusion methods in remote sensing.
@@
Initially, convolution neural networks (CNNs) were the dominant architecture, known for their ability to extract local spatial features from different modalities with the encoder-decoder framework [UNet, li2021abcnet]. These early CNN-based methods stack multimodal data and fuse features at various stages for improved segmentation performance [FuseNet, zhang2022multilevel]. With the introduction of the self-attention-based Transformer [tf], which excels at modeling global context and long-range dependencies, many hybrid methods integrating Transformers into CNN-based methods have emerged. In particular, Vision Transformer (ViT) [vit] and Swin Transformer [liu2021swin] further introduced Transformer to the computer vision community, greatly improving the ability of image feature extraction. Combining CNNs for detailed feature extraction with Transformers for capturing global relationships between different data sources marked a new stage in segmentation models [ma2022crossmodal, he2023mftransnet, ma2024frequency]. This hybrid approach enhances the ability to fuse information across modalities and scales, leading to more accurate, robust segmentation results in complex scenes. Despite their good performance, the aforementioned models were trained solely on narrowly scoped task-specific data, limiting their acquisition of general visual knowledge. In contrast, foundation models benefit from large-scale, diverse pretraining data, enabling them to develop broad visual representations that extend beyond single-task constraints.
Figure 1: Illustration of the key challenges: Traditional task-specific remote sensing models are limited in general visual knowledge. In contrast, SAM contains general knowledge from large-scale natural image corpora, but it lacks the framework to support multimodal remote sensing tasks.
@@
Interestingly, the recent emergence of large foundation models offers a solution to this challenge. In particular, SAM [kirillov2023segment] is a cutting-edge segmentation model designed to tackle a wide range of segmentation tasks across diverse datasets. It is comprised of three components, including a ViT-based image encoder, a prompt encoder and a mask decoder. Developed by Meta AI, SAM benefits from large-scale training on a vast visual corpus of natural images, enabling it to generalize effectively to unseen objects. Its versatility and robust performance position it as a valuable tool for various applications [ma2024segment, wang2024sam]. However, compared to natural images, remote sensing images exhibit significant differences due to variations in sensors, resolution, spectral ranges and so on [li2018deep, zheng2020foreground, ma2023unsupervised]. For multimodal tasks, non-optical data such as DSM further increases the discrepancy. This raises a new challenge: How can the capabilities of SAM, acquired from massive natural visual corpora, be leveraged to enhance multimodal remote sensing tasks? Fig. 1 illustrates the key challenges discussed above, providing a clearer depiction of the problem addressed in this work.
@@
Existing fine-tuning techniques, especially Adapter [houlsby2019parameter, chen2022adaptformer, he2023parameter] and Low-Rank Adaptation (LoRA) [hu2021lora, zhang2023adaptive], address part of this challenge. Compared to training and full fine-tuning, these methods fix most of the parameters and learn task-specific information with very few parameters. This approach enables parameter-efficient learning and the successful migration of large foundation models to a wider range of specific downstream tasks, even in more constrained hardware environments. The concept of Adapter was first introduced in the natural language processing community [houlsby2019parameter] as a method to fine-tune large pretrained models for specific downstream tasks. The core idea behind Adapter is to introduce a parallel, compact, and scalable adaptation module that learns task-specific knowledge during training while the original model branch remains fixed, retaining task-agnostic knowledge. Similarly, LoRA [hu2021lora] introduced trainable rank decomposition matrices to learn task-specific knowledge. This approach leverages the synergy between task-specific and task-agnostic knowledge, enabling efficient fine-tuning of the large model.
Figure 2: (a) Current multimodal fusion methods, and (b) The proposed unified multimodal fine-tuning framework. In our method, single-modal encoders are used to learn general knowledge and frozen, enabling multimodal tasks to benefit from the vision foundational model.
@@
In the field of remote sensing, fine-tuning techniques have focused on adapting SAM for single-modality tasks [pu2024classwise, zhou2024mesam, lu2024multi]. For instance, CWSAM [pu2024classwise] and MeSAM [zhou2024mesam] adapted SAM’s image encoder and introduced custom mask decoders for remote sensing data. However, an analysis of the SAM parameter distribution across three scales reveals that the majority of SAM’s parameters are concentrated in its image encoder, suggesting that most of its general knowledge is encapsulated within this component. Therefore, we consider it unnecessary to modify SAM’s prompt encoder and mask decoder for remote sensing-specific tasks, as such modifications increase model adaptation complexity while hindering seamless integration with existing segmentation models.
To address the aforementioned challenges with fine-tuning framework, we present a unified multimodal fine-tuning framework for multimodal data feature learning and fusion. Specifically, as illustrated in Fig. 2(a), current multimodal fusion methods typically assign separate encoders for each modality and perform feature fusion through a multimodal fusion module [ResUNet-a, ma2024ftransunet, zhang2024asanet]. When additional modalities are introduced, more encoders need to be added accordingly. During training, all encoders and the fusion module need to be trained simultaneously. In contrast, the unified multimodal fine-tuning framework proposed in this work revolutionizes this structure. As depicted in Fig. 2(b), it capitalizes on the general knowledge embedded in the foundational model and facilitates modality learning and fusion through the multimodal fine-tuning module. During training, the encoders remain fixed, requiring optimization only for the multimodal fine-tuning module. This approach not only fully leverages the vision foundational model but also offers seamless scalability to incorporate additional modalities.
@@
After that, to thoroughly investigate the impact of various fine-tuning mechanisms in remote sensing, we introduce two multimodal fine-tuning techniques, namely Multimodal Adapter (MMAdapter) and Multimodal LoRA (MMLoRA), for multimodal fusion remote sensing tasks. It is worth noting that our multimodal fine-tuning framework represents a novel feature learning fusion strategy that operates independently of the specific fine-tuning module used. Leveraging MMAdapter and MMLoRA, respectively, we propose a novel multimodal fusion method, namely Multimodal Fine-tuning Network (MFNet) with different image encoder architectures. Furthermore, MFNet utilizes a Deep Fusion Module (DFM) to perform multiscale processing and fusion of the deep high-level features from SAM’s image encoder. This is an effective solution for the complex characteristics of remote sensing data. Additionally, we employ a universal semantic segmentation decoder [wang2022unetformer] that does not require additional task-specific design efforts, allowing for easy integration with decoders from other tasks. The fourfold contributions are summarized in the following:
  • A unified multimodal fine-tuning framework is proposed to adapt SAM for multimodal feature learning, independent of the specific design of the fine-tuning module and the number of data modalities;
@@
  • A scalable multimodal fusion network, namely MFNet, is proposed by capitalizing on SAM’s image encoder and the proposed multimodal fine-tuning framework to perform remote sensing semantic segmentation. This is a streamlined and flexible adaptation network that eliminates most of the redundant modules in SAM;
@@
  • Two most representative fine-tuning architectures from the literature, namely Adapter and LoRA, are utilized to validate the framework’s effectiveness by showing SAM’s robust generalization capability on DSM for the first time. Extensive experiments on three well-known multimodal remote sensing datasets, ISPRS Vaihingen, ISPRS Potsdam and MMHunan, confirm that the proposed MFNet substantially outperforms existing methods in terms of semantic segmentation performance;
  • To the best of our knowledge, as pioneers in exploring the multimodal fine-tuning approach based on SAM, this work thoroughly investigates the role of the two most widely used fine-tuning mechanisms in the field of remote sensing. It establishes a robust foundation for related studies and provides clear directions for future research.
@@
The remainder of this paper is structured as follows. Sec. II provides a review of related works on multimodal fusion and SAM. Sec. III presents the unified multimodal fine-tuning framework and MFNet, followed by a detailed description and analysis of the extensive experiments and a discussion in Sec. IV. Finally, the conclusion is given in Sec. V.
@@ II-A Multimodal Remote Sensing Semantic Segmentation
Semantic segmentation is a critical preprocessing step in remote sensing image processing, and leveraging multimodal information often yields better results than relying on a single modality. Recently, the advent of deep learning has revolutionized the entire field of remote sensing, including semantic segmentation. Based on the classical encoder-decoder framework [UNet, li2021abcnet], numerous multimodal fusion approaches based on CNNs and Transformers have driven significant advancements in the field [vfusenet, ma2022crossmodal, he2023mftransnet, ma2024ftransunet]. ResUNet-a [ResUNet-a], an early CNN-based architecture, simply stacked multimodal data into four channels. Furthermore, vFuseNet [vfusenet] introduced a two-branch encoder to separately extract multimodal features, enabling deeper multimodal fusion through element-wise operations at the feature level. Recently, the introduction of Transformers [tf, vit] has further enriched multimodal networks. For instance, CMFNet [ma2022crossmodal] employed CNNs for feature extraction and uses a Transformer structure to connect multimodal features across scales, emphasizing the importance of scale in multimodal fusion. Similarly, MFTransNet [ma2024ftransunet] used CNNs for feature extraction, while enhancing the self-attention module with spatial and channel attention for better feature fusion. FTransUNet [ma2024ftransunet] presented a multilevel fusion approach to refine the fusion of shallow- and deep-level remote sensing semantic features. Despite the great performance they achieved, we argue that existing models lack sufficient general knowledge, which poses a fundamental limitation to the advancement of multimodal fusion methods.
Figure 3: (a) The ViT block without Adapter in SAM’s image encoder, (b) the ViT block equipped with standard Adapter [wu2023medical], (c) the ViT block endowed with the proposed MMAdapter and (d) the detailed structure of the MMAdapter. The Adapter facilitates the efficient utilization of general knowledge in specific tasks. Compared to standard Adapter, MMAadapter is characterized by two branches of shared weights for multimodal feature extraction. The standard Adapter and the proposed MMAdapter are used to fine-tune and fuse features, respectively.
@@ II-B SAM in Remote Sensing
The SAM [kirillov2023segment] holds a unique position as a general-purpose image segmentation model. This vision foundation model was trained on a very large visual corpus. It endows SAM with a remarkable ability to generalize to unseen objects, making it well-suited for applications in diverse scenarios. Nowadays, SAM has already been applied across various fields, such as autonomous driving [li2024fusionsam], medical image processing [mazurowski2023segment] and remote sensing [SAMRS, qi2024samp, chen2024change, ma2024sam]. In remote sensing, SAMRS [SAMRS] leveraged SAM to integrate numerous existing remote sensing datasets, using a new prompt named rotated bounding box. Furthermore, some recent works have considered fine-tuning SAM for remote sensing tasks such as semantic segmentation [pu2024classwise, zhou2024mesam, yan2023ringmo] and change detection [ding2023adapting, mei2024scd].
For single modality tasks, CWSAM [pu2024classwise] adapted SAM to synthetic aperture radar (SAR) images by introducing a task-specific input module and a class-wise mask decoder. MeSAM [zhou2024mesam] incorporated an inception mixer into the SAM’s image encoder to retain high-frequency features and introduced a multiscale-connected mask decoder for optical images. SAM_MLoRA [lu2024multi] employed multiple LoRA modules in parallel to enhance LoRA’s decomposition capability. For multimodal tasks, RingMo-SAM [yan2023ringmo] introduced a prompt encoder tailored to multimodal remote sensing data, along with a category-decoupling mask decoder. It is observed that these methods focus on refining the fine-tuning mechanisms and designing task-specific prompts or mask decoders. They have preliminarily explored the generalization ability of SAM in remote sensing tasks. However, as discussed in Sec. I, the general knowledge in SAM is primarily centered in the image encoder. While these approaches successfully utilize SAM’s general knowledge to remote sensing, their complex architectures greatly hinder their adaptation to existing general semantic segmentation networks. Furthermore, there is currently no SAM-based multimodal approach designed for DSM data.
@@ III Proposed Method
We first introduce the unified multimodal fine-tuning framework by elaborating the MMAdapter and MMLoRA. Specifically, we review the conventional single-modal fine-tuning strategy Adapter and the proposed MMAdapter in Sec. III-A. After that, we present another classic single-modal fine-tuning strategy LoRA and the proposed MMLoRA in Sec. III-B. Building upon the proposed multimodal fine-tuning mechanisms, the proposed MFNet is elaborated in Sec. III-C. Notably, MFNet has two distinct architectures based on the selection between MMAdapter and MMLoRA. Finally, to provide a clear explanation, we use the example of two modalities for illustration.
III-A Standard Adapter and the Proposed MMAdapter
  • Standard Adapter: As discussed in Sec. I, the general knowledge is confined to SAM’s image encoders, specifically the ViT blocks, whose structure is shown in Fig.3(a). In [wu2023medical], the Adapter was proposed to enhance the capabilities of the ViT blocks for medical tasks through fine-tuning, as illustrated in Fig. 3(b). Instead of adjusting all parameters, the pretrained SAM parameters remain frozen, while two Adapter modules are introduced to learn task-specific knowledge. Each Adapter consists of a down projection, a ReLU activation, and an up projection. The down projection compresses the input embedding to a lower dimension using a simple MLP layer, and the up projection restores the compressed embedding to its original dimension using another MLP layer. For a specific input feature $\bm{x}_{i}\in\mathbb{R}^{h\times w\times c}$, where $h$, $w$, and $c$ represent the height, width, and channels of the input feature, the Adapter’s process for generating the adapted feature can be expressed as:
    $\bm{x}^{SA}_{a}$ $=\mathrm{ReLU}\left(\mathrm{LN}(\bm{x}_{i})\cdot\bm{W}_{d}\right)\cdot\bm{W}_{u},$ (1)